MICROWAVE SPECTROSCOPIC STUDY OF THE CO-H_{2}S COMPLEX

\underline{YUNJIE XU}, PATRICIA ARBOLEDA, and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada, T6G 2G2.

The infrared spectra of the van der Waals complex between CO and H_{2}S were recently studied by Xia and McKellar.^{a} Following this investigation, we have measured rotational spectra of several isotopomers of CO-H_{2}S using a cavity Fourier transform microwave spectrometer. Nuclear hyperfine structures due to the quadrupolar ^{33}S and ^{17}O nuclei were observed and analyzed. The resulting rotational and hyperfine constants were used to derive structural and dynamical information about the complex.

aC. Xia and A. R. W. McKellar, private communication